OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=4,(6,2)

OpenStudy (katrinakaif):

SAME AS BEFORE! :) Plug it into the equation (y - y1) = m (x -x1) M being the slope and with your ordered pair (0,7) as (x1, y1)

OpenStudy (katrinakaif):

But this time your ordered pair is (6 2)

OpenStudy (anonymous):

The ordered pair is (6,2) this time though, Otherwise the process is the same

OpenStudy (anonymous):

Tried it, cant figure it out. This one seems more complicated

OpenStudy (katrinakaif):

Okay So first you know the slope is 4 (m) (y -y1) = 4( x - x1) Now plug in the ordered pairs (6, 2) as (x1,y1) Gives us the euqation: (y - 2) = 4( x -6) Now can you solve?

OpenStudy (anonymous):

next step would be to distribute the 4. Try that and let us know what you get.

OpenStudy (anonymous):

Distribute the four? How do I do that?

OpenStudy (anonymous):

\[a(b - c) = a(b) - a(c)\]

OpenStudy (anonymous):

so in this case: a is 4 b is x c is 6